# Language Models can Self-Improve at State-Value Estimation for Better Search

### v2

@@ Abstract
Collecting ground-truth rewards or human demonstrations for multi-step reasoning tasks is often prohibitively expensive and time consuming, especially in interactive domains like web tasks. To address this bottleneck, we present self-taught lookahead (STL), a self-supervised method that leverages state-transition dynamics to improve a value model capable of effectively guiding language model-controlled search without any labeled data. We find that moderately sized (8 billion parameters) open-weight value models improved with STL can match the performance of using a gpt-4o value model. Furthermore, we find that specialized value models learned with STL can be deployed with computationally lightweight search algorithms, achieving performance that matches that of more expensive tree search methods, while reducing costs by an order of magnitude.
@@ 1 Introduction
While large language models (LLMs) demonstrate strong reasoning capabilities by generating extended token sequences before answering wei2022chain ; guo2025deepseek ; muennighoff2025s1 , guiding inference with explicit tree search has the potential to further improve performance on tasks with a structured state space yao2024tree ; xie2024self . In this setting, a policy LLM proposes candidate actions, and a value LLM evaluates resulting states to steer the search toward promising trajectories. Figure 1 summarizes the assumptions different LLM-driven search methods make about information available during training and inference. Reward and Demo Learning methods yao2022webshop ; zhang2024rest_mcts ; jin2025search ; song2025r1 assume access to ground-truth reward signals or human demonstrations and optimize the model with reinforcement learning (RL) or imitation learning (IL). In contrast, Reward-Guided Inference strategies zhou2023lats ; shinn2024reflexion forego explicit reward during training; they only consult a reward signal at inference time, using it to guide procedures such as LLM-based Monte-Carlo Tree Search (MCTS). However, collecting ground-truth rewards or human demonstrations may not be possible in every environment, and can oftentimes be costly. For example, for web agent tasks, even small-scale data collection can cost thousands of dollars yao2022webshop . Reward and Demo Free methods yao2022react ; yao2024tree relax this assumption as they can operate without access to reward. However, these methods often rely on prompting an off-the-shelf LLM to serve as both the policy and value models during the search process, which constrains performance to the capabilities of the underlying LLM.
@@
By representing the lookahead process in natural language rather than regressing solely on value estimates, STL takes advantage of LLMs’ strong generalization to unseen tasks via learned textual reasoning nye2021show ; zelikman2022star . For instance, our results (§4) on web agent tasks from WebShop yao2022webshop demonstrate that tree search with an STL fine-tuned llama-3.1-8b-instruct value model improves performance by $39\%$ or more compared to the base llama model. Furthermore, STL matches the performance of search with a base gpt-4o value model and even achieves comparable results to Reward-Guided Inference methods such as LATS zhou2023lats on unseen tasks. We show that these results also hold for math puzzles and multi-hop question answering yao2024tree ; yang-etal-2018-hotpotqa . Finally, by enabling the use of a small open-source value model during inference, STL leads to significant cost reduction when agents are deployed. Through an efficiency analysis (§5), we find that STL costs 5$\times$ less than a similarly performing gpt-4o value model. As STL produces more accurate state-value estimates, the resulting value models can effectively guide search algorithms that expand fewer states, while maintaining strong performance.
@@ 3 Better State-Value Estimation with Self-Taught Lookahead
In this section, we present our proposed self-taught lookahead method. §3.1 and §3.2 describe the STL improvement procedure, while §3.3 explains how a STL-improved value model operates during inference. See Figure 2 and Algorithm 1 in Appendix A for an overview of the method.
@@ 3.1 Generating Rollouts
where $\gamma$ is the discount factor. Since tasks are episodic, we set $\gamma=1$ bernstein2001reinforcement ; van2005survey ; leffler2007efficient . These lookahead values capture a better estimate of the true value of $s_{j}$ as they account for $s_{j}$’s successor states. In §6, we describe how generating and learning from these $y$’s is similar to fitted value iteration.
@@
$o_{s_{j}}\leftarrow a^{*}_{j}||s^{*}_{j+1}||\text{Rat}(V_{\phi_{k}}(x_{i},s_{0},\ldots,s_{j},s^{*}_{j+1}))$ (4)
where |$\cdot$|$\cdot$ | | ⋅ denotes concatenation, $s^{*}_{j+1}\leftarrow T(s_{j},a^{*}_{j})$, and
$a^{*}_{j}\in\operatorname*{arg\,max}_{a\in A_{s_{j}}}\Big{\{}\text{Val}(V_{\phi_{k}}(x_{i},s_{0},\ldots,s_{j},T(s_{j},a)))\Big{\}}\$ (5)
@@ 4.1 Web Tasks
[…]
@@
Reflexion gpt-3.5-turbo —— 58.5 24.0 49.5 16.4
[…]
@@
Following the empirical advantages on agent tasks identified by yu2024exact , we generate training data with MCTS by performing a step of lookahead at each step during rollout. Note that we use the LLM value model value outputs as a proxy reward to guide UCT (Upper Confidence bounds applied to Trees) kocsis2006bandit selection like yu2024exact , instead of ground truth reward zhou2023lats . We perform STL with a gpt-3.5-turbo gpt-35 policy to be consistent with previous work zhou2023lats ; shinn2024reflexion as well as a gpt-4o GPT-4o policy and fine-tune a llama-3.1-8b-instruct dubey2024llama value function. STL is performed by rolling out 50 tasks from the WebShop training set, resulting in close to 1500 training examples, which we find is sufficient for significant performance improvement. Also, while we perform data generation during STL with MCTS, we evaluate the agent using the trained value models with greedy search, where the next action is greedily chosen based on the values of the policy’s proposed actions. Finally, we find that a single iteration of STL is sufficient to see significant improvement over using a base LLM-initialized value model, and that multiple iterations do not yield additional performance improvements due to difficulties in simulating more than one step ahead, given the complexity of the environment. This conclusion has been corroborated by other LLM agent works gu2024your ; chae2024web . However, in §4.3 we show that this multi-step simulation is possible for simpler tasks. See Appendix D for further details about data generation, training, and simulation.
@@
Within Reward and Demo Learning approaches, we include IL and RL methods originally proposed in the WebShop work that train BERT devlin2019bert and BART lewis2019bart models on human demonstrations and ground truth reward yao2022webshop . We additionally compare to the current state-of-the-art approach AgentQ putta2024agent , which finetunes a larger xLAM-v0.1-r-46.7b policy on rolled out MCTS search trees using direct preference optimization rafailov2023direct . We also include Reward-Guided Inference methods such as Reflexion shinn2024reflexion and LATS zhou2023lats , which work by prompting closed-source LLMs within a framework, e.g., MCTS, that is guided by ground truth reward. Finally, in the Reward and Demo Free setting, we include greedy search and MCTS [^2] We use LLM value as a proxy reward to guide UCT, like in the data generation phase of STL baselines with a ReACT yao2022react prompted base LLM (llama-3.1-8b-instruct, gpt-3.5-turbo, gpt-4o, r1-distill-llama-8b guo2025deepseek ). All value models are prompted with few-shot examples and are asked to provide reasoning before a numerical value. For clear comparisons, a branching factor of 5 is used for greedy and MCTS-based methods like LATS, i.e., LATS is strictly more computationally expensive than greedy methods. We use the pass@$3$ chen2021passk for methods that do not have access to reward at inference time.
@@
The WebShop average reward (Score) and success rate (SR) of all methods evaluated are presented in Table 1. We present results on both the full WebShop test set and on the mini test set of 50 tasks used by zhou2023lats , as we find running LATS and other MCTS methods on the entire test set is computationally expensive. Both of these sets are distinct from those seen during STL. Within the Reward and Demo Free setting, we find that STL matches the performance of using a gpt-4o value model and leads to a greater than 7% improvement in average reward and a 39% improvement in success rate relative to a base llama-3.1-8b-instruct value model, both of which are statistically significant improvements ($p<0.001$) using the paired bootstrap test berg2012empirical . Moreover, we find that STL even outperforms Reward-Guided Inference methods that have access to ground truth rewards. Lastly, we find no statistically significant difference between using a gpt-3.5-turbo or gpt-4o policy due to high action diversity in our setup (see Appendix D.2); thus, we evaluate search methods with a single policy (either gpt-3.5-turbo or gpt-4o) in subsequent experiments.
Table 2: Ablation study on the impact of fine-tuning with different combinations of information from lookahead namely lookahead values (lv), textual representation of the next best action and successor state (tr), and the value rationale for the successor state (r). The underlined results are from the base model before any fine-tuning.
@@
In the Reward and Demo Free setting, we also perform ablations on the set of information from the step of lookahead used to fine-tune the value model during self-improvement. Specifically, we compare STL to variants that use only subsets of the information derived from lookahead. As mentioned in §3, this information includes (1) the lookahead value, (2) the textual representation of the next-best action and its successor state, and (3) the value rationale for the successor state. The results of this ablation in Table 2 demonstrate that regressing solely on lookahead values and further incorporating state transitions from lookahead does improve performance relative to the base model. However, learning also from the value rationale of the successor state, as done in STL, yields additional performance gains over these other settings. These results substantiate the claims made in §3 about the necessity of learning from action-outcome rationales, a key difference between STL and both classical RL gordon1999fitted_value_iteration and other LLM tree search works feng2023alphazero ; zhang2024rest_mcts which fine-tune an LLM value model on numerical values only.
@@ 4.2 Multi-Hop Question Answering
[…]
@@
[…]
@@ 4.3 Math Puzzles
Figure 3 shows the performance of evaluated methods on a set of 50 tasks seen during STL and a set of 50 more challenging (determined by lower human solve percentages), unseen tasks. On both sets, STL matches or outperforms a gpt-4o value model. However, STL’s performance on seen tasks monotonically increases for the first three iterations, while its performance on unseen tasks decreases before increasing. This phenomenon may be due to the base value model generalizing better than self-improved models in earlier iterations, during which the model has not received signal from more accurate LLM values of terminal states during lookahead.
@@ 5.1 Compute and Cost Efficiency
Keeping compute requirements and costs low is critical, especially for agents automating routine, repetitive tasks. Since STL can be used to improve an open-source LLM like llama-3.1-8b-instruct, we can transfer computation from more expensive closed-source models like gpt-4o to open-source alternatives while maintaining performance. Figure 4 (right) demonstrates this transfer, as STL uses more than 50% fewer closed-source tokens than greedy search with a gpt-3.5-turbo or gpt-4o value model. We also compute the monetary cost of different methods, accounting for both inference costs and, for STL, those incurred during fine-tuning. See Appendix G for details about the cost calculation. We plot these costs against the average WebShop reward in Figure 5 (left), and find that STL is Pareto optimal, 23×23$\times$23 × cheaper than MCTS methods like LATS, and 5×5$\times$5 × cheaper than performing greedy search with a gpt-4o value model.
Figure 5: Tradeoff between performance and efficiency on WebShop with a gpt-3.5-turbo policy. Pareto frontiers of existing methods and baselines are shown, illustrating the optimality of STL when considering the tradeoff between cost and average reward (left) and between environmental usage and average reward (right). Fine-tuning costs are included in the total STL cost.
@@ 6 Related Work
The majority of prior work on training reasoning agents focuses on performing SFT on human-annotated trajectories yao2022webshop ; lai2024autowebglm , synthetically generated trajectories chen2023fireact ; furuta2023multimodal ; zhang2024xlam ; liu2024apigen ; patel2024large ; murty2024nnetscape ; ou2024synatra ; su2025learn , or a combination of both zeng2023agenttuning ; zhang2024agentohana . Other work has trained agents from tree search generated data, but these methods usually require ground truth reward gandhi2024stream ; putta2024agent ; zhang2024rest_mcts . While prior work has explored self-improving reasoning agents, these approaches fail to generalize beyond the instructions encountered during self-improvement patel2024large or require fine-tuning frontier models like gpt-4o to achieve generalization yu2024exact .

### v3

@@ Abstract
Collecting ground-truth rewards or human demonstrations for multi-step reasoning tasks is often prohibitively expensive, particularly in interactive domains such as web tasks. We introduce Self-Taught Lookahead (STL), a reward-free framework that improves language model–based value functions by reasoning explicitly about state transitions. STL can be viewed as a chain-of-thought analogue of the value iteration algorithm: instead of regressing directly on numeric values, a value LLM is trained to simulate a step of lookahead in natural language—predicting the next action, resulting state, and rationale for its value, thereby refining value estimates without any labeled data. This self-supervised procedure yields more accurate state-value predictions, which in turn enable lightweight search algorithms to expand fewer states while maintaining strong performance. Empirically, STL-trained value models built on moderately sized (8B parameter) open-weight LLMs boost web agent success rates by 39%, achieving comparable performance with proprietary models. STL also generalizes to multi-hop QA and math puzzles. We find that STL enables small open-source models to guide efficient search, reducing inference costs by integrating explicit reasoning with value learning.
@@ 1 Introduction
While large language models (LLMs) demonstrate strong reasoning capabilities by generating extended token sequences before answering wei2022chain ; guo2025deepseek ; muennighoff2025s1 , guiding inference with explicit tree search has the potential to further improve performance on tasks with a structured state space yao2024tree ; xie2024self . In this setting, a policy LLM proposes candidate actions, and a value LLM evaluates resulting states to steer the search toward promising trajectories. Figure 1 summarizes the assumptions different LLM-driven search methods make about information available during training and inference. Reward and Demo Learning methods yao2022webshop ; zhang2024rest_mcts ; jin2025search ; song2025r1 assume access to ground-truth reward signals or human demonstrations and optimize the model with reinforcement learning (RL) or imitation learning (IL). In contrast, Reward-Guided Inference strategies zhou2023lats ; shinn2024reflexion forego explicit reward during training; they only consult a reward signal at inference time, using it to guide procedures such as LLM-based Monte-Carlo Tree Search (MCTS). However, collecting ground-truth rewards or human demonstrations may not be possible in every environment, and can oftentimes be costly. For example, for web agent tasks, even small-scale data collection can cost thousands of dollars yao2022webshop . Reward and Demo Free methods yao2022react ; yao2024tree relax this assumption as they can operate without access to reward. However, these methods often rely on prompting an off-the-shelf LLM to serve as both the policy and value models during the search process, which constrains performance compared to models specifically tuned for agentic tasks wei2025browsecomp .
@@
[…]
@@ 3 Better State-Value Estimation with Self-Taught Lookahead
In this section, we present our proposed self-taught lookahead method. §3.1 and §3.2 describe the STL improvement procedure, while §3.3 explains how a STL-improved value model operates during inference. See Figure 2 and Algorithm 1 in Appendix C for an overview of the method.
@@ 3.1 Generating Rollouts
where $\gamma$ is the discount factor. Since tasks are episodic, we set $\gamma=1$ bernstein2001reinforcement ; van2005survey ; leffler2007efficient . These lookahead values capture a better estimate of the true value of $s_{j}$ as they account for $_{j}’s s$uccessor states. In §7, we describe how generating and learning from these $y$’s is similar to fitted value iteration.
@@
[…]
where $\cdot||\cdot$ denotes concatenation, $s^{*}_{j+1}\leftarrow T(s_{j},a^{*}_{j})$, and
$a^{*}_{j}\in\operatorname*{arg\,max}_{a\in A_{s_{j}}}\Big\{\text{Val}(V_{\phi_{k}}(x_{i},s_{0},\ldots,s_{j},T(s_{j},a)))\Big\}\\$ (5)
@@ 4.1 Web Tasks
Table 1: Score and success rate (SR) on WebShop. Results marked with $\dagger$}} are taken from previous work zhou2023lats ; putta2024agent . Value functions marked with $\ddagger$} are fine-tuned. We observe a near 40% improvement in success rate when using the STL value function compared to the llama-3.1-8b-instruct base value model in the greedy setting. We compute statistical significance of Reward and Demo Free methods against the underlined results ($^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001)$ using the paired bootstrap test berg2012empirical . Best results in the Reward and Demo Free setting are bolded.
@@
—— —— —— —— 50.5
Reflexion gpt-3.5-turbo —— 77.2 46.0 72.9 41.3
Reward-Guided Inference LATS gpt-3.5-turbo gpt-3.5-turbo 75.9 38.0 —— ——
@@
Following the empirical advantages on agent tasks identified by yu2024exact , we generate training data with MCTS by performing a step of lookahead at each step during rollout. Note that we use the LLM value model value outputs as a proxy reward to guide UCT (Upper Confidence bounds applied to Trees) kocsis2006bandit selection like yu2024exact , instead of ground truth reward zhou2023lats . We perform STL with a gpt-3.5-turbo gpt-35 policy to be consistent with previous work zhou2023lats ; shinn2024reflexion as well as a gpt-4o GPT-4o policy and fine-tune a llama-3.1-8b-instruct dubey2024llama value function. STL is performed by rolling out 50 tasks from the WebShop training set, resulting in 1161 training examples, which we find is sufficient for significant performance improvement. We find that training a separate value model at each depth allows us to train with smaller LLMs with fewer active parameters during rollouts (see Appendix D.3 for more details). Also, while we perform data generation during STL with MCTS, we evaluate the agent using the trained value models with greedy search, where the next action is greedily chosen based on the values of the policy’s proposed actions (see Algorithm 2 for more details). Finally, we find that a single iteration of STL is sufficient to see significant improvement over using a base LLM-initialized value model, and that multiple iterations do not yield additional performance improvements due to difficulties in simulating more than one step ahead, given the complexity of the environment. This conclusion has been corroborated by other LLM agent works gu2024your ; chae2024web . However, in §4.3 we show that this multi-step simulation is possible for simpler tasks. See Appendix D for further details about data generation, training, and simulation.
@@
Within Reward and Demo Learning approaches, we include IL and RL methods originally proposed in the WebShop work that train BERT devlin2019bert and BART lewis2019bart models on human demonstrations and ground truth reward yao2022webshop . We additionally compare to the current state-of-the-art approach AgentQ putta2024agent , which finetunes a larger xLAM-v0.1-r-46.7b policy on rolled out MCTS search trees using direct preference optimization rafailov2023direct . We also include Reward-Guided Inference methods such as Reflexion shinn2024reflexion and LATS zhou2023lats , which work by prompting closed-source LLMs within a framework, e.g., MCTS, that is guided by ground truth reward. Finally, in the Reward and Demo Free setting, we include greedy search and MCTS [^2] We use LLM value as a proxy reward to guide UCT, like in the data generation phase of STL baselines with a ReACT yao2022react prompted base LLM (llama-3.1-8b-instruct, gpt-3.5-turbo, gpt-4o, r1-distill-llama-8b guo2025deepseek ). All value models are prompted with few-shot examples and are asked to provide reasoning before a numerical value. Following LATS for a fair comparison, we use a branching factor of 5 for all methods and 30 iterations for MCTS-based approaches. This makes LATS and MCTS baselines strictly more computationally expensive than greedy methods. We use the pass@$3$ chen2021passk for methods that do not have access to reward at inference time. Finally, we include a comparison to human expert performance measured in the original WebShop paper yao2022webshop .
@@
The WebShop average reward (Score) and success rate (SR) of evaluated methods are presented in Table 1, with additional baselines in Table 4. We present results on both the full WebShop test set and on the mini test set of 50 tasks used by zhou2023lats , as we find running LATS and other MCTS methods on the entire test set is computationally expensive. Both of these sets are distinct from those seen during STL. Within the Reward and Demo Free setting, we find that STL matches the performance of using a gpt-4o value model and leads to a greater than 7% improvement in average reward and a 39% improvement in success rate relative to a base llama-3.1-8b-instruct value model, both of which are statistically significant improvements ($p<0.001$) using the paired bootstrap test berg2012empirical . Moreover, we find that STL even performs similarly to Reward-Guided Inference methods that have access to ground truth rewards. Lastly, we find no statistically significant difference between using a gpt-3.5-turbo or gpt-4o policy due to high action diversity in our setup (see Appendix D.2); thus, we evaluate search methods with a single policy (either gpt-3.5-turbo or gpt-4o) in subsequent experiments.
Table 2: Ablation study on the impact of fine-tuning with different combinations of information from lookahead, namely lookahead values (lv), textual representation of the next best action and successor state (tr), and the value rationale for the successor state (r). The underlined results are from the base model before any fine-tuning.
@@
In the Reward and Demo Free setting, we also perform ablations on the set of information from the step of lookahead used to fine-tune the value model during self-improvement. Specifically, we compare STL to variants that use only subsets of the information derived from lookahead. As mentioned in §3, this information includes (1) the lookahead value, (2) the textual representation of the next-best action and its successor state, and (3) the value rationale for the successor state. The results of this ablation in Table 2 demonstrate that regressing solely on lookahead values and further incorporating state transitions from lookahead does improve performance relative to the base model. However, learning also from the value rationale of the successor state, as done in STL, yields additional performance gains over these other settings. These results substantiate the claims made in §3 about the necessity of learning from action-outcome rationales, a key difference between STL and both classical RL gordon1999fitted_value_iteration and other LLM tree search works feng2023alphazero ; zhang2024rest_mcts which fine-tune an LLM value model on numerical values only. Note for fairness, we modify the loss-masking to improve the lookahead value-only baseline and include this higher result in the table. See Appendix F.4 for more details about loss masking and this baseline.
@@ 4.2 Multi-Hop Question Answering
Table 3: Match rates on HotpotQA. Value functions marked with $\ddagger$}} are fine-tuned. Statistical significance with the paired bootstrap test of Reward and Demo Free methods against the underlined results ($^{***}p<0.001$) is provided. The best  Reward and Demo Free results are bolded.
@@
Reward-Guided Inference LATS gpt-3.5-turbo gpt-3.5-turbo 70.0
@@ 4.3 Math Puzzles
Figure 3 shows the performance of evaluated methods on a set of 50 tasks seen during STL and a set of 50 more challenging (determined by lower human solve percentages), unseen tasks. On both sets, STL matches or outperforms a gpt-4o value model. However, STL’s performance on seen tasks monotonically increases for the first three iterations, while its performance on unseen tasks decreases before increasing. This phenomenon is due to the limited number of tasks the value model sees during training during the first couple of iterations. Specifically, if value models are not exposed to enough actions and their lookahead values during training, they fail to generalize well to unseen tasks. Limiting the number of tasks per iteration also limits the quantity and diversity of actions and values seen. A full analysis of this result can be found in Appendix F.5.
@@ 5.1 Compute and Cost Efficiency
Keeping compute requirements and costs low is critical, especially for agents automating routine, repetitive tasks. Since STL can be used to improve an open-source LLM like llama-3.1-8b-instruct, we can transfer computation from more expensive closed-source models like gpt-4o to open-source alternatives while maintaining performance. Figure 4 (right) demonstrates this transfer, as STL uses more than 50% fewer tokens generated from closed-source than greedy search with a gpt-3.5-turbo or gpt-4o value model. We also compute the monetary costs of different methods. Unlike other methods, STL incurs costs for data generation and fine-tuning in addition to inference, but these are one-time costs that do not scale with agent use and are quite modest at \$8.54 and \$1.76, respectively. We plot inference costs against the average WebShop reward in Figure 5 (left), and find that STL is Pareto optimal, $23\times$ cheaper than MCTS methods like LATS, and $5\times$ cheaper than performing greedy search with a gpt-4o value model. See Appendix G for details about the cost calculation.
Figure 5: Tradeoff between performance and efficiency on WebShop with a gpt-3.5-turbo policy. Pareto frontiers of existing methods and baselines are shown, illustrating the optimality of STL when considering the tradeoff between inference cost and average reward (left) and between environmental usage and average reward (right). Reward-Guided Inference methods are presented in gold and not included in the Pareto frontier since they belong to a different information setting.
@@ 5.3 STL Scaling Trends
6 Discussion: When to use STL?
As discussed in §4 and further detailed in Appendix D.2, achieving strong performance with STL requires high action diversity (a large number of possible actions at each step). When action diversity is low, the effect of the value model on search performance is diminished, and it is necessary to roll out more tasks to obtain enough data to fine-tune the value model. For example, Section 4.2 notes that due to its low action diversity, HotpotQA shows smaller relative gains from STL compared to WebShop and also requires rollouts of 500 tasks (10 times more than WebShop) to get a large enough dataset for fine-tuning.
Additionally, STL requires later states to provide good value estimates that can be backed up and learned during training on lookahead results. Search performance will benefit most from STL on tasks where state transitions are consistent throughout the environment, i.e., the same or semantically similar actions yield similar outcomes. For instance, clicking the “Low to High” button on any search results page consistently orders items by price. This consistency enables the STL value model to accurately simulate a step of lookahead (§3.3), leading to better state-value estimation and improved downstream search performance. Tasks that have stochastic transitions or have inconsistent transitions when actions are semantically similar may not show large improvements with STL. Fortunately, most popular tasks where LLM agents have been deployed, such as web navigation, have deterministic and fairly consistent transitions.
@@ 7 Related Work
The majority of prior work on training reasoning agents focuses on performing SFT on human-annotated trajectories yao2022webshop ; lai2024autowebglm , synthetically generated trajectories chen2023fireact ; furuta2023multimodal ; zhang2024xlam ; liu2024apigen ; patel2024large ; murty2024nnetscape ; ou2024synatra ; su2025learn , or a combination of both zeng2023agenttuning ; zhang2024agentohana . Other work has trained agents from tree search-generated data gandhi2024stream ; putta2024agent ; zhang2024rest_mcts or with explicit reinforcement learning  peiyuan2024agile ; wei2025webagent ; feng2025group , but these methods usually require ground truth reward. While prior work has explored self-improving reasoning agents, these approaches fail to generalize beyond the instructions encountered during self-improvement patel2024large or require fine-tuning frontier models like gpt-4o to achieve generalization yu2024exact .